OpenStudy (acespeedfighter):

34 + 4 ⋅ 5 =____ . (Input only whole numbers.) (1 point) I think its 105 check please

jimthompson5910 (jim_thompson5910):

how are you getting 105? can you post your steps please?

OpenStudy (acespeedfighter):

@jim_thompson5910 why do you have a high level 99 but you dont have a PFP

jimthompson5910 (jim_thompson5910):

haven't gotten around to it I guess

jimthompson5910 (jim_thompson5910):

anyways, how did you get 105?

OpenStudy (acespeedfighter):

4*5=20

OpenStudy (acespeedfighter):

3*3*3*3=81+4 = 85

OpenStudy (acespeedfighter):

85+20=105

jimthompson5910 (jim_thompson5910):

oh i see, it's 3^4 and not 34

jimthompson5910 (jim_thompson5910):

so the original expression is this? \[\Large 3^4+4*5\]

OpenStudy (acespeedfighter):

yes

OpenStudy (acespeedfighter):

i forgot to do 3^4

jimthompson5910 (jim_thompson5910):

3^4 = 3*3*3*3 = 81 4*5 = 20 add up the two results: 81+20 = 101

OpenStudy (acespeedfighter):

what about ohhhh i see my mistake thx for cheching it i used 4 2 times

jimthompson5910 (jim_thompson5910):

no problem